OpenStudy (tinylittlehelper):

Help please! Kareena has $20. She woke up and went shopping. She spent money on 5 apples, and each is $2. She drove to work and payed for 3 gallons on 46 cent a gallon. On the way, she found a $5 bill on the ground. When Kareena arrived, she worked PART of her night shift, which consists of $10 an hour and had 8 hours in all. Then she worked over time for 3 hours and earned 6/47 of what she earned on her night shift. She worked for 8 hours in total. She paid for dinner for $4 and 7.5% tax included. At the end of her day, how much money did she have?

OpenStudy (anonymous):

hang on lol

OpenStudy (anonymous):

nevermind i g2g but just if she bought $10 worth of apples and $1.38 on gas and then get the tax on $4 then add it then subtract all that from what she earned total

OpenStudy (tinylittlehelper):

Is this expression right for the problem? Anyone? 20 -(10 + (0.46 x 3)) + 5 = A A + ((8 - 3) x 10) = B B + ((6/47) x B) - (4 + (4 x 7.5%)) = ????

OpenStudy (tinylittlehelper):

Anyone?